3. A hydraulic press has a 30mm ratio input piston and 75mm radio output piston. What input force is needed to provide an output force that is able to lift a truck 1300kg?
1
Expert's answer
2018-09-06T16:53:08-0400
P_1=P_2
F_1/A_1 =F_2/A_2
F_2=F_1 (A_2/A_1 )=F_1 (r_2/r_1 )^2
We have
F_1=mg.
Thus,
F_2=mg(r_2/r_1 )^2=(1300)(9.8) (30/75)^2=2040 N.